• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 16:12:02 on August25,2006

    Entry number 182515

    keyword=Day Shift Summary

    8:00 AM Day Shift Begins
    Shift Workers Lindgren(SL), Calarco(3rd), adn Glister(TO)
    Change kinematics LHRS P0 to 0.3080, Change magnet currents
    Increase beam current to 20 ua
    Change to LD2 Target
    8:37 AM Start Run 3292 LD2 500K Events Production Run
    9:11 AM Start Run 3293 LD2 500K Events Production Run
    9:45 AM Start Run 3294 LD2 500K Events Production Run
    10:25 AM LHRS P0 change to 0.3573 and magnet current change
    !0:30 AM Lost bean due to IOC software problem
    11:01 AM till waiting for beam
    11:30 AM Beam back 20.7 ua, chnage PS factors
    11:30 Start Run 3295 4 cm Dumy target P0=0.3573
    11:34 Stop Run 3295 PS not set right, Raise PS factor RHRS
    11:38 Start Run 3296

    11:34 Stop Run 3295 PS not set right, Raise PS factor RHRS
    11:38 Start Run 329611:30 AM Beam back 20.7 ua, chnage PS factors
    11:30 Start Run 3295 4 cm Dumy target P0=0.3573
    11:34 Stop Run 3295 PS not set right, Raise PS factor RHRS
    11:38 Start Run 329611:30 AM Beam back 20.7 ua, chnage PS factors
    11:30 Start Run 3295 4 cm Dumy target P0=0.3573
    11:34 Stop Run 3295 PS not set right, Raise PS factor RHRS
    12:30 Beam Back 20.9 ua
    12:30 Start Run 3297 4 cm Dummy 250K Events
    12:40 PM Comment: At the end of the runs for momentum setting 0.3573
    go back to momentum settings 0.3080 and 0.2681 and take 4 cm dummy cell runs.
    13:05 Stop Run 3297
    Change Target to Carbon, Set PS factors
    13:10 Start Run 3298 Carbon Target 100K Events
    13:41 Start Run 3299 Carbon Target 100K Events
    14:07 Start Run 3300 Carbon Target 100K Events
    14:40 Start Run 3301 Carbon Raster Off 100K Events
    Change Target to Ta, Turn raster on, Reset PS factors
    15:15 Start Run Ta target, raster on, Spot is okay
    15:45 Stop Run Ta target 100K Events
    Left HRS Gaussmeter stuck, Several reboots did not help,need to make Hall A access
    16:00 Shift Ends